# Symmetric square type L–series

### v1

@@ 5.2. Rankin Selberg realisation
In 1939 Rankin [Rankin1939II] considered integrals of automorphic functions multiplied by a non-holomorphic Eisenstein series of weight zero over the fundamental domain of $\operatorname{SL}_{2}(\mathbb{Z})$ given by
@@
which represents a fundamental domain of $\Gamma_{\infty}\leq\operatorname{SL}_{2}(\mathbb{Z})$ being the stabiliser of $\infty$. The resulting integral would represent an $L$-series with analytic continuation and functional equation [Rankin1939II, Thm. 3 p. 360]. We remark that Rankin already considered the case of Hecke congruence subgroups and that the method was independently developed by Selberg.
@@
  • $f\in\mathcal{S}_{L,k}$ with Fourier coefficients $a(\lambda,n)\in\mathcal{O}_{\varepsilon}(n^{\nu_{f}})$ for some $\R\nu_{f}\in\R$,
  • $g\in\mathcal{M}_{L_{1},k_{1}}$ with Fourier coefficients $b(\lambda,n)\in\mathcal{O}_{\varepsilon}(n^{\nu_{g}})$ for some $\R\nu_{g}\in\R$,
@@
    The substitution $v\mapsto v\cdot N/[2\pi(\lvert n\rvert+\lvert n_{1}\rvert)]$ yields that the last expression equals
@@
[…]
@@
=\;= $\int_{0}^{\infty}\sum_{\lambda_{1}\in L_{1}^{\prime}/L_{1}}\sum_{0\neq l\in L_{1}+\lambda_{1}}a(\lambda_{1}+\eta,\operatorname{q}(l))\cdot e[iv2\operatorname{q}(l)]\cdot v^{\overline{s}+k-2}\;\textnormal{d}v$
=\;= $\frac{\Gamma(\overline{s}+k-1)}{(4\pi)^{\overline{s}+k-1}}\sum_{0\neq l\in L_{1}^{\prime}}\frac{a(\overline{l}+\eta,\operatorname{q}(l))}{\operatorname{q}(l)^{\overline{s}+k-1}}.$
@@ 5.3. Product expansions
  • Proof:
    Recall that the Fourier coefficients of $f\in\mathcal{S}_{\mathcal{L},k}$ are denoted $a(\lambda,n)$ as in (3.1). Following [ShimuraMF_HalfIntegralWeight, p. 452], define for $n\in\mathbb{N}$ the formal power series
@@
    =\;= $\sum_{\begin{subarray}{c}n=1\\ \gcd(n,Np)\end{subarray}}^{\infty}H_{n}(p^{-s})n^{-s}=\;L_{(\lambda,t)}^{Np}(f,s)\cdot Q(p^{-s}).$
@@
A key property of absolutely convergent products is that such a product attains the value zero, if, and only if, at least one of its factors vanishes. We refer the reader to [Freitag1, IV p. 200] for more on that matter and proceed by applying Proposition 5.8 in order to derive infinite product expansions.
@@
    If we prove that $P_{p}(p^{-s})$ converges to $1$, then it suffices to verify that $R_{p}(p^{-s},t)-P_{p}(p^{-s})$ converges to 00 faster than $1/p$. As before, we distinguish the odd and even case, however, before specialising to one of these, recall that by Lemma 4.11 the following bound is true:
@@
    $\left\lvert P_{p}(p^{-s})-1\right\rvert\leq\;$ $\left\lvert\frac{\sigma_{p}}{p^{k-1}}-(1-p^{-1})\frac{G_{\mathcal{L}}(1)}{G_{\mathcal{L}}(p)}\right\rvert p^{k-1-\Re(s)}+p^{2(k-1-\Re(s))}$
    =\;= $\left(\frac{p+1}{p}+\frac{1-p^{-1}}{p}\right)\cdot p^{k-\Re(s)}+p^{2(k-1-\Re(s))}.$
    We see that for $\Re(s)>k$, the expression converges to 00 for $p\to\infty$ as desired. Next, the numerator is investigated:
@@
    As a consequence, the difference $R_{p}(p^{-s},t)-P_{p}(p^{-s})$ goes to zero for $p\to\infty$ in case of $\Re(s)>k$. It goes to 00 faster than $1/p$ for $\Re(s)>k+1$. This settles absolute convergence of the product in the even case.
@@
    $\lvert P_{p}(p^{-s})-1\rvert\leq\;$ $\left\lvert\frac{\sigma_{p}}{p^{k-1}}\right\rvert\cdot p^{k-1-\Re(s)}+p^{2(k-1-\Re(s))}$
    =\;= $\frac{p+1}{p}\cdot p^{k-\Re(s)}+p^{2(k-1-\Re(s))}.$
    We see that for $\Re(s)>k$, the expression converges to 00 for $p\to\infty$. Next, the numerator is investigated:
@@
    As a consequence, the difference $Q_{p}(p^{-s},t)-P_{p}(p^{-s})$ goes to zero for $p\to\infty$ in case $\Re(s)>k$. It goes to 00 faster than $1/p$ for $\Re(s)>k+1$. This settles absolute convergence of the product in the odd case.
@@
    $\sigma_{p}a(\lambda,n)=\;$ $p^{2(k-1)}\delta_{\lambda\in{{}^{p}\!\mathcal{L}}}\sum_{\begin{subarray}{c}\lambda^{\prime}\in{{}_{p}\mathcal{L}}\\ n-p^{2}\operatorname{\overline{q}}(\lambda/p+\lambda^{\prime})\in p^{2}\mathbb{Z}\end{subarray}}a\left(\lambda/p+\lambda^{\prime},\frac{n-p^{2}\operatorname{\overline{q}}(\lambda/p+\lambda^{\prime})}{p^{2}}+\operatorname{\overline{q}}(\lambda/p+\lambda^{\prime})\right)$
    +\;+ $p^{k-2}\delta_{\lambda\in{{}^{p}\!\mathcal{L}}}K_{L,p}g_{p}\left[1,{{}_{-}\chi_{p}}^{R_{p}},n-\operatorname{q}(p\cdot\ell_{\lambda/p})\right]\;a\left(\lambda,n\right)$
    +\;+ $a(p\lambda,p^{2}n).$
    This results for some $n,m\in\mathbb{N}$ with $p\nmid n$ and $p^{2}\nmid\operatorname{lev}(L)t$ or $\lambda_{p}\neq 0$ in111 Note that the group $\mathcal{L}_{p}$ may only contain constituents of the form $\mathbb{Z}/p\mathbb{Z}$.
@@
    +\;+ $p^{k-2}K_{L,p}g_{p}\left[1,{{}_{-}\chi_{p}}^{R_{p}},(p^{m}n)^{2}t-\operatorname{q}(p\cdot\ell_{(p^{m}n\lambda)/p}))\right]a\!\left(p^{m}n\lambda,(p^{m}n)^{2}t\right)$
    (II) +\;+ $a(p^{m+1}n\lambda,(p^{m+1}n)^{2}t).$
@@
    We deduce that the Gauss sum is indifferent to multiplication of the last argument by a square that is coprime to $p$. Multiplication by $(\operatorname{lev}(L)/p^{\nu_{p}(\operatorname{lev}(L))})^{2}$ yields that the number $(\operatorname{lev}(L)/p^{\nu_{p}(\operatorname{lev}(L))})^{2}p^{2}\operatorname{q}(\ell_{\lambda/p})$ is an integer divisible by $p$ [^2] Recall that we have assumed $\mathcal{L}_{p}$ to be non-isotropic, implying $p^{2}\nmid\operatorname{lev}(L)$. Hence, $\delta_{p\nmid t-\operatorname{q}(p\ell_{\lambda/p})}=\delta_{p\nmid t}$ and this expression is already implicitly included in the Legendre symbol, yielding
@@
    Should we have $\lambda_{p}\neq 0$, only the case $m=1$ remains. Then the separation in (5.9) works when replacing $\lambda$ by $\lambda-\lambda_{p}$, shifting the non-trivial $p$-component to $\lambda^{\prime}$ this time. Note that there will be at least two choices of $\lambda^{\prime}\in{{}_{p}\mathcal{L}}$ then. Namely, $\lambda_{p}\neq 0$ and $-\lambda_{p}$. In fact, there might be even more and the number of these choices $\lambda^{\prime}\in\mathcal{L}_{p}$ will be denoted by $C(\lambda_{p})+1$. To advance the computation, the different Fourier coefficients appearing in the sum in (II) have to be related.333 Note that the sum in equation (II) introduces indices corresponding to various choices of $\lambda^{\prime}$ that were not present in the original series $L_{(\lambda,t)}^{N}(f,s)$. This is where the assumption of invariance with respect to the action of the orthogonal group comes into play. Recall that for $\lambda_{p}\neq 0$, we find $\operatorname{\overline{q}}(\lambda_{p})\neq 0$ by assumption on the lattice. This implies the maximality of the local lattice $L_{p}$ which renders $\mathcal{L}_{p}$ an $\mathbb{F}_{p}$ vector space. Then, mapping $\lambda_{p}$ to a different choice of root with the same norm defines an isometry of subspaces, which, by Witts extension theorem, extends to an isometry of $\mathcal{L}_{p}$, i.e. an element in $\operatorname{O}(\mathcal{L}_{p})$. By assumption, the Fourier coefficients of $f$ are invariant with respect to this action so that we obtain $C(\lambda_{p})+1$ contributions from this particular coefficient and identify $C(\lambda_{p})+1$ as the size of the orbit of $\lambda_{p}$ under $\operatorname{O}(\mathcal{L}_{p})$.
@@
    +\;+ $\delta_{2\mid{R_{p}}}p^{k-2}K_{L,p}\cdot(p-1)\;a\left(p^{m}n\lambda,(p^{m}n)^{2}t\right)$
    (II’) +\;+ $a(p^{m+1}n\lambda,(p^{m+1}n)^{2}t).$
@@
    =\;= $\delta_{\lambda_{p}\neq 0}(p-1)-\delta_{\lambda_{p}=0}p(\delta_{p\mid t}-1)$
    =\;= $\delta_{\lambda_{p}\neq 0}(p-1)+\delta_{\lambda_{p}=0}\delta_{p\nmid t}p.$
@@
    =\;= $a(n\lambda,n^{2}t)\times\left[1+p^{k-2}K_{L,p}\begin{cases}\delta_{\lambda_{p}\neq 0}(p-1)+\delta_{\lambda_{p}=0}\delta_{p\nmid t}p,&2\mid{R_{p}}\\ -\delta_{\lambda_{p}=0}p^{1/2}\left(\frac{-t}{p}\right)\epsilon_{p},&2\nmid{R_{p}}\end{cases}\cdot x-C(\lambda_{p})p^{2(k-1)}x^{2}\right]$
@@
    =\;= $a(n\lambda,n^{2}t)\cdot\frac{1+K_{L,p}\begin{cases}\delta_{\lambda_{p}\neq 0}(1-p^{-1})+\delta_{\lambda_{p}=0}\delta_{p\nmid t},&2\mid{R_{p}}\\ -\delta_{\lambda_{p}=0}p^{-1/2}\left(\frac{-t}{p}\right)\epsilon_{p},&2\nmid{R_{p}}\end{cases}\cdot p^{k-1-s}-C(\lambda_{p})p^{2(k-1-s)}}{1-\left(\sigma_{p}-\delta_{2\mid{R_{p}}}(1-p^{-1})p^{k-1}K_{L,p}\right)p^{-s}+p^{2(k-1-s)}}.$
@@
    in this right half plane. Note that in the above equation no factor vanishes, unless $R(p^{-s},t)$ does. To this end, we determine a right half plane in which $|1-R(p^{-s},t)\rvert<1$. The first factor that needs to be tended to is $K_{L,p}$. By [ScheithauerWeilRep, Prop. 3.8] we find
@@
[…]
=\;= $\begin{cases}\sum_{n\in\mathbb{N}}\frac{a(2^{n}\lambda,2^{n}t)}{2^{ns}},&2\mid\operatorname{lev}(L),\\ a(\lambda,t),&2\nmid\operatorname{lev}(L)\end{cases}$
@@
[…]

### v2

@@ 5.2. Rankin Selberg realisation
In 1939 Rankin [Ran39] considered integrals of automorphic functions multiplied by a non-holomorphic Eisenstein series of weight zero over the fundamental domain of $\operatorname{SL}_{2}(\mathbb{Z})$ given by
@@
which represents a fundamental domain of $\Gamma_{\infty}\leq\operatorname{SL}_{2}(\mathbb{Z})$ being the stabiliser of $\infty$. The resulting integral would represent an $L$-series with analytic continuation and functional equation [Ran39, Thm. 3 p. 360]. We remark that Rankin already considered the case of Hecke congruence subgroups and that the method was independently developed by Selberg.
@@
  • $f\in\mathcal{S}_{L,k}$ with Fourier coefficients $a(\lambda,n)\in\mathcal{O}_{\varepsilon}(n^{\nu_{f}})$ for some $\nu_{f}\in\mathbb{R}$,
  • $g\in\mathcal{M}_{L_{1},k_{1}}$ with Fourier coefficients $b(\lambda,n)\in\mathcal{O}_{\varepsilon}(n^{\nu_{g}})$ for some $\nu_{g}\in\mathbb{R}$,
@@
[…]
@@
Clearly, $n-\operatorname{q}(l)\in\mathbb{Z}$, so that the mapping $\mathbb{R}\ni u\mapsto e\left[u(n-\operatorname{q}(l))\right]$ defines a character which descends to a character on the unit circle $\mathbb{T}$. This implies
@@
$=\;$ $\int_{0}^{\infty}\sum_{\lambda_{1}\in L_{1}^{\prime}/L_{1}}\sum_{0\neq l\in L_{1}+\lambda_{1}}a(\lambda_{1}+\eta,\operatorname{q}(l))\cdot e[iv2\operatorname{q}(l)]\cdot v^{\overline{s}+k-2}\;\textnormal{d}v$
$=\;$ $\frac{\Gamma(\overline{s}+k-1)}{(4\pi)^{\overline{s}+k-1}}\sum_{0\neq l\in L_{1}^{\prime}}\frac{a(\overline{l}+\eta,\operatorname{q}(l))}{\operatorname{q}(l)^{\overline{s}+k-1}}.$
@@ 5.3. Product expansions
Recall that the Fourier coefficients of $f\in\mathcal{S}_{\mathcal{L},k}$ are denoted $a(\lambda,n)$ as in (3.1). Following [Shi73, p. 452], define for $n\in\mathbb{N}$ the formal power series
@@
$=\;$ $\sum_{\begin{subarray}{c}n=1\\ \gcd(n,Np)\end{subarray}}^{\infty}H_{n}(p^{-s})n^{-s}=\;L_{(\lambda,t)}^{Np}(f,s)\cdot Q(p^{-s}).$
@@
A key property of absolutely convergent products is that such a product attains the value zero, if, and only if, at least one of its factors vanishes. We refer the reader to [FB06, IV p. 200] for more on that matter and proceed by applying Proposition 5.8 in order to derive infinite product expansions.
@@
If we prove that $P_{p}(p^{-s})$ converges to $1$, then it suffices to verify that $R_{p}(p^{-s},t)-P_{p}(p^{-s})$ converges to 0 faster than $1/p$. As before, we distinguish the odd and even case, however, before specialising to one of these, recall that by Lemma 4.11 the following bound is true:
@@
[…]
$=\;$ $\left(\frac{p+1}{p}+\frac{1-p^{-1}}{p}\right)\cdot p^{k-\Re(s)}+p^{2(k-1-\Re(s))}.$
We see that for $\Re(s)>k$, the expression converges to 0 for $p\to\infty$ as desired. Next, the numerator is investigated:
@@
As a consequence, the difference $R_{p}(p^{-s},t)-P_{p}(p^{-s})$ goes to zero for $p\to\infty$ in case of $\Re(s)>k$. It goes to 0 faster than $1/p$ for $\Re(s)>k+1$. This settles absolute convergence of the product in the even case.
@@
[…]
$=\;$ $\frac{p+1}{p}\cdot p^{k-\Re(s)}+p^{2(k-1-\Re(s))}.$
We see that for $\Re(s)>k$, the expression converges to 0 for $p\to\infty$. Next, the numerator is investigated:
@@
As a consequence, the difference $Q_{p}(p^{-s},t)-P_{p}(p^{-s})$ goes to zero for $p\to\infty$ in case $\Re(s)>k$. It goes to 0 faster than $1/p$ for $\Re(s)>k+1$. This settles absolute convergence of the product in the odd case.
@@
$\sigma_{p}a(\lambda,n)=\;$ $p^{2(k-1)}\delta_{\lambda\in{{}^{p}\!\mathcal{L}}}\hskip-5.0pt\sum_{\begin{subarray}{c}\lambda^{\prime}\in{{}_{p}\mathcal{L}}\\ n-p^{2}\operatorname{\overline{q}}(\lambda/p+\lambda^{\prime})\in p^{2}\mathbb{Z}\end{subarray}}\hskip-5.0pta\left(\lambda/p+\lambda^{\prime},\frac{n-p^{2}\operatorname{\overline{q}}(\lambda/p+\lambda^{\prime})}{p^{2}}+\operatorname{\overline{q}}(\lambda/p+\lambda^{\prime})\right)$
$+\;$ $p^{k-2}\delta_{\lambda\in{{}^{p}\!\mathcal{L}}}K_{L,p}g_{p}\left[1,{{}_{-}\chi_{p}}^{R_{p}},n-\operatorname{q}(p\cdot\ell_{\lambda/p})\right]\;a\left(\lambda,n\right)$
$+\;$ $a(p\lambda,p^{2}n).$
This results for some $n,m\in\mathbb{N}$ with $p\nmid n$ and $p^{2}\nmid\operatorname{lev}(L)t$ or $\lambda_{p}\neq 0$ in [^1] Note that the group $\mathcal{L}_{p}$ may only contain constituents of the form $\mathbb{Z}/p\mathbb{Z}$.
@@
$+\;$ $p^{k-2}K_{L,p}g_{p}\left[1,{{}_{-}\chi_{p}}^{R_{p}},(p^{m}n)^{2}t-\operatorname{q}(p\cdot\ell_{(p^{m}n\lambda)/p}))\right]a\!\left(p^{m}n\lambda,(p^{m}n)^{2}t\right)$
(II) $+\;$ $a(p^{m+1}n\lambda,(p^{m+1}n)^{2}t).$
@@
We deduce that the Gauss sum is indifferent to multiplication of the last argument by a square that is coprime to $p$. Multiplication by $(\operatorname{lev}(L)/p^{\nu_{p}(\operatorname{lev}(L))})^{2}$ yields that the number $(\operatorname{lev}(L)/p^{\nu_{p}(\operatorname{lev}(L))})^{2}p^{2}\operatorname{q}(\ell_{\lambda/p})$ is an integer divisible by pp. [^2] Recall that we have assumed $\mathcal{L}_{p}$ to be non-isotropic, implying $p^{2}\nmid\operatorname{lev}(L)$. Hence, $\delta_{p\nmid t-\operatorname{q}(p\ell_{\lambda/p})}=\delta_{p\nmid t}$ and this expression is already implicitly included in the Legendre symbol, yielding
@@
Should we have $\lambda_{p}\neq 0$, only the case $m=1$ remains. Then the separation in (5.9) works when replacing $\lambda$ by $\lambda-\lambda_{p}$, shifting the non-trivial $p$-component to $\lambda^{\prime}$ this time. Note that there will be at least two choices of $\lambda^{\prime}\in{{}_{p}\mathcal{L}}$ then. Namely, $\lambda_{p}\neq 0$ and $\lambda_{p}$. In fact, there might be even more and the number of these choices $\lambda^{\prime}\in\mathcal{L}_{p}$ will be denoted by $C(\lambda_{p})+1$. To advance the computation, the different Fourier coefficients appearing in the sum in (II) have to be related. [^3] Note that the sum in equation (II) introduces indices corresponding to various choices of $\lambda^{\prime}$ that were not present in the original series $L_{(\lambda,t)}^{N}(f,s)$. This is where the assumption of invariance with respect to the action of the orthogonal group comes into play. Recall that for $\lambda_{p}\neq 0$, we find $\operatorname{\overline{q}}(\lambda_{p})\neq 0$ by assumption on the lattice. This implies the maximality of the local lattice $L_{p}$ which renders $\mathcal{L}_{p}$ an $\mathbb{F}_{p}$ vector space. Then, mapping $\lambda_{p}$ to a different choice of root with the same norm defines an isometry of subspaces, which, by Witts extension theorem, extends to an isometry of $\mathcal{L}_{p}$, i.e. an element in $\operatorname{O}(\mathcal{L}_{p})$. By assumption, the Fourier coefficients of $f$ are invariant with respect to this action so that we obtain $C(\lambda_{p})+1$ contributions from this particular coefficient and identify $C(\lambda_{p})+1$ as the size of the orbit of $\lambda_{p}$ under $\operatorname{O}(\mathcal{L}_{p})$.
@@
$+\;$ $\delta_{2\mid{R_{p}}}p^{k-2}K_{L,p}\cdot(p-1)\;a\left(p^{m}n\lambda,(p^{m}n)^{2}t\right)$
(II’) $+\;$ $a(p^{m+1}n\lambda,(p^{m+1}n)^{2}t).$
@@
$=\;$ $\delta_{\lambda_{p}\neq 0}(p-1)-\delta_{\lambda_{p}=0}p(\delta_{p\mid t}-1)$
$=\;$ $\delta_{\lambda_{p}\neq 0}(p-1)+\delta_{\lambda_{p}=0}\delta_{p\nmid t}p.$
@@
$=\;$ $a(n\lambda,n^{2}t)\times\left[1+p^{k-2}K_{L,p}\begin{cases}\delta_{\lambda_{p}\neq 0}(p-1)+\delta_{\lambda_{p}=0}\delta_{p\nmid t}p,&2\mid{R_{p}}\\ -\delta_{\lambda_{p}=0}p^{1/2}\left(\frac{-t}{p}\right)\epsilon_{p},&2\nmid{R_{p}}\end{cases}\cdot x-C(\lambda_{p})p^{2(k-1)}x^{2}\right]$
@@
$=\;$ $a(n\lambda,n^{2}t)\cdot\frac{1+K_{L,p}\begin{cases}\delta_{\lambda_{p}\neq 0}(1-p^{-1})+\delta_{\lambda_{p}=0}\delta_{p\nmid t},&2\mid{R_{p}}\\ -\delta_{\lambda_{p}=0}p^{-1/2}\left(\frac{-t}{p}\right)\epsilon_{p},&2\nmid{R_{p}}\end{cases}\cdot p^{k-1-s}-C(\lambda_{p})p^{2(k-1-s)}}{1-\left(\sigma_{p}-\delta_{2\mid{R_{p}}}(1-p^{-1})p^{k-1}K_{L,p}\right)p^{-s}+p^{2(k-1-s)}}.$
@@
in this right half plane. Note that in the above equation no factor vanishes, unless $R(p^{-s},t)$ does. To this end, we determine a right half plane in which $|1-R(p^{-s},t)\rvert<1$. The first factor that needs to be tended to is $K_{L,p}$. By [Sch09, Prop. 3.8] we find
@@
$\overset{\textnormal{def}}{=}\;$ $\sum_{n\in\mathbb{N}}\frac{a(n\lambda,n^{2}t)}{n^{s}}$
$=\;$ $\begin{cases}\sum_{n\in\mathbb{N}}\frac{a(2^{n}\lambda,2^{n}t)}{2^{ns}},&2\mid\operatorname{lev}(L),\\ a(\lambda,t),&2\nmid\operatorname{lev}(L)\end{cases}$
@@
Do note that in the range of absolute convergence specified above, the $L$-series vanishes, if, and only if, $\sum_{n\in\mathbb{N}}a(2^{n}\lambda,2^{n}t)/2^{ns}=0$ in case $2\mid\operatorname{lev}(L)$ or $a(\lambda,t)=0$ in case $2\nmid\operatorname{lev}(L)$.